# Privacy-Accuracy Trade-offs in High-Dimensional LASSO under Perturbation Mechanisms

### 27 Mar 2026 v1

@@ 1 Introduction
A common approach to achieve differential privacy is to randomize the process of parameter estimation, a principle known as the randomized mechanism. Intuitively, such randomization prevents an adversary from reliably inferring whether a particular individual’s data was used in the analysis, by ensuring that the algorithm’s output remains probabilistically similar even when one data point is modified. Several randomized mechanisms have been proposed for statistical estimation. For instance, (Dwork et al., 2006) introduced the Laplace mechanism, in which independent Laplace noise is added to the output of a query (i.e., a function of the dataset, such as an estimator) in proportion to its sensitivity, providing a concrete method to achieve $(\epsilon,0)$-differential privacy. This method belongs to the so-called output perturbation framework, where noise is added directly to the estimated model parameters after optimization. Another widely used approach is objective perturbation (or loss perturbation), where noise is injected into the loss function before or during optimization (Chaudhuri et al., 2011). These two classes of mechanisms exhibit different theoretical and empirical behaviors in terms of privacy guarantees and statistical accuracy, although in specific cases, such as least-squares estimation, they can be shown to be mathematically equivalent under suitable conditions.
A fundamental challenge in differential privacy lies in balancing the trade-off between preserving privacy and maintaining the usefulness of data, often referred to as the privacy-utility trade-off (Dwork et al., 2014; Chaudhuri et al., 2011). Utility in this context includes statistical accuracy, model interpretability, computational feasibility, and application-specific relevance. This trade-off is intrinsic to privacy-preserving mechanisms: stronger privacy guarantees typically require introducing more randomness, which can deteriorate the utility of the resulting model or statistical output. In supervised learning tasks, a prominent instance of the privacy-utility trade-off is the privacy-accuracy trade-off. Here, utility is measured in terms of predictive accuracy or estimation error, and the goal is to quantify how privacy mechanisms degrade the model’s performance.
Output perturbation is often simpler to analyze, but it may fail to adapt to structural properties of the underlying parameters. In particular, in sparse estimation problems such as LASSO, adding noise directly to the estimated coefficients can destroy sparsity patterns and deteriorate the interpretability and statistical efficiency of the solution (Khanna et al., 2023). To address this issue, several works have proposed variants of output perturbation that explicitly account for sparsity. A common approach is to first estimate the support of the parameters and then apply output perturbation only to the selected (nonzero) components (Lei et al., 2018). Similar ideas appear in differentially private model selection procedures, where variable selection is performed prior to private estimation on the reduced model (Smith, 2011; Talwar et al., 2015). These approaches can improve utility by avoiding unnecessary noise on irrelevant coordinates, and have been shown to achieve favorable accuracy-privacy trade-offs under sparsity assumptions. However, such two-stage procedures introduce additional challenges from a privacy perspective. Since the support selection step itself depends on the data, it may leak sensitive information unless it is properly privatized. By the composition property of differential privacy (Dwork et al., 2014), this typically requires allocating a portion of the privacy budget to the model selection step, which increases the overall complexity of the mechanism and may degrade performance. In contrast, objective perturbation injects noise directly into the loss function before optimization. This approach implicitly accounts for the regularization structure of the estimator, and can better preserve properties such as sparsity without requiring an explicit support selection step. This motivates us to analyze objective perturbation as an alternative mechanism within a unified framework.
In the context of statistical estimation, particularly penalized regression methods such as the LASSO, several works have developed perturbation-based mechanisms to achieve differential privacy guarantees. However, these analyses typically rely on strong structural assumptions on the data or the design matrix, such as restricted eigenvalue conditions or exact sparsity (Chaudhuri et al., 2011; Talwar et al., 2015; Smith, 2011), which may not hold in practical high-dimensional settings. Existing results suggest that objective perturbation can outperform output perturbation under strong convexity and smoothness assumptions (Chaudhuri et al., 2011). However, it remains unclear whether such advantages persist in high-dimensional sparse regimes, where the estimator is inherently non-smooth. Moreover, the interplay between privacy noise, regularization, and generalization performance remains only partially understood.
@@
To better understand these interactions, we focus on the high-dimensional regime, where both the number of samples and parameters grow to infinity with a fixed ratio. In such settings, worst-case guarantees from differential privacy can be overly conservative, as they are often dominated by rare pathological configurations that do not reflect typical data realizations. In contrast, typical-case analysis captures the behavior under random design and noise, which is more relevant for modern statistical applications. Despite its importance, this perspective has received limited attention, and in particular, the role of sparsity in shaping effective privacy leakage has not been systematically characterized.
In such settings, typical-case analysis requires a distributional characterization of the estimator under random data. This naturally motivates an information-theoretic approach to quantifying privacy leakage. Specifically, we consider the problem of distinguishing whether an estimator was computed from a dataset $\mathcal{D}$ or from a neighboring dataset $\mathcal{D}^{\prime}$ differing in a single sample. This naturally leads to a binary hypothesis testing problem, where the optimal error exponent is governed by the KL divergence between the corresponding distributions. This connection provides an operational interpretation of KL-based privacy measures: a small KL divergence implies that even an optimal adversary cannot reliably infer the presence of an individual data point.
@@
Building on this formulation, we analyze the high-dimensional LASSO under both output perturbation and objective perturbation using tools from approximate inference, in particular approximate message passing (AMP). AMP provides an asymptotic characterization of high-dimensional estimators under random design, enabling a typical-case analysis of both generalization error and privacy-related quantities. This framework allows us to quantify how sparsity, noise strength, and the sample-to-parameter ratio jointly influence both prediction accuracy and privacy, without relying on restrictive conditions such as restricted eigenvalue assumptions or worst-case sensitivity bounds.
@@ 1.1 Related Works
A substantial body of work has developed differentially private variants of regularized regression methods, including ridge regression and LASSO, using techniques including output perturbation, objective perturbation, and noise injection into sufficient statistics (Chaudhuri et al., 2011; Talwar et al., 2015). While these methods provide rigorous privacy guarantees under certain modeling assumptions, their theoretical guarantees in high-dimensional settings often rely on structural conditions-such as restricted eigenvalue assumptions, which can be difficult to satisfy in overparameterized regimes.
In the broader context of private empirical risk minimization, Bassily et al. (2014) established minimax optimal rates, clarifying the fundamental trade-off between privacy and utility. While these worst-case guarantees are foundational, they can be overly conservative and may not reflect the typical behavior of learning algorithms. Related developments, including proximal perturbation methods for non-smooth optimization (Wang et al., 2018) and compressed noise injection techniques (Zhou et al., 2020), further expand the scope of private optimization, but remain primarily focused on worst-case analyses.
Complementary to these works, on-average privacy measures such as the On-average KL divergence (OnAveKL) (Wang et al., 2016) have been proposed to capture typical-case privacy leakage. These metrics admit an operational interpretation in terms of hypothesis testing and membership inference, and provide a more refined view of privacy in practice. However, existing studies have largely focused on low-dimensional or unregularized settings, leaving open the question of how high dimensionality, sparsity, and regularization jointly affect on-average privacy.
Despite these advances, a systematic understanding of privacy-accuracy trade-offs in high-dimensional sparse models remains lacking. In particular, the interaction between noise injection mechanisms, structural constraints such as sparsity, and typical-case privacy measures has not been fully characterized.
@@ 1.2 Our contribution
  • Implications for noise calibration: Our analytical characterization provides practical guidance for selecting the noise level to balance prediction accuracy and privacy. In particular, we identify regimes where increasing noise leads to non-monotonic effects on stability and privacy under objective perturbation.
Taken together, our results provide a unified and quantitative understanding of privacy-accuracy trade-offs in high-dimensional sparse estimation, complementing existing worst-case analyses.
@@ 7 Conclusion
$\mathrm{m}_{i}^{(t)}$ $={\Sigma}_{i}^{(t)}\sum_{\mu=1}^{n}\left\{X_{\mu i}\frac{y_{\mu}-\widehat{y}_{\mu}^{\setminus\mu}}{1+{s_{\theta}}_{\mu}}+\frac{X_{\mu i}^{2}\widehat{\beta}_{i}^{(t)}}{1+{s_{\theta}}_{\mu}}\right\}$ (72)
$\frac{1}{{\Sigma}_{i}^{(t)}}$ $=\sum_{\mu=1}^{n}\frac{X^{2}_{\mu i}}{1+{s_{\theta}}}.$ (73)
@@
$=\frac{1}{\sum_{\mu=1}^{n}X^{2}_{\mu i}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{i=1}^{p}X_{\mu i}\beta_{i}^{(0)}+\xi_{\mu}-\sum_{i=1}^{p}X_{\mu i}\widehat{\beta}_{i\to\mu})+X_{\mu i}^{2}\widehat{\beta}_{i\to\mu}^{(t)}\right\}$
$=\frac{1}{\sum_{\mu=1}^{n}X^{2}_{\mu i}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{i=1}^{p}X_{\mu i}\beta_{i}^{(0)}+\xi_{\mu})-\sum_{j\neq i}^{p}X_{\mu i}X_{\mu j}\widehat{\beta}_{j\to\mu}\right\}$ (74)
@@
$\mathrm{m}_{i}^{(t)}-\beta_{i}^{(0)}$ $=\frac{1}{\sum_{\mu=1}^{n}X^{2}_{\mu i}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{j=1}^{p}X_{\mu j}\beta_{j}^{(0)}+\xi_{\mu})-\sum_{j\neq i}^{p}X_{\mu i}X_{\mu j}\widehat{\beta}_{j\to\mu}-X_{\mu i}^{2}\beta_{i}^{(0)}\right\}$
$=\frac{1}{\sum_{\mu=1}^{n}X^{2}_{\mu i}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{j\neq i}^{p}X_{\mu j}\beta_{j}^{(0)}+\xi_{\mu})-\sum_{j\neq i}^{p}X_{\mu i}X_{\mu j}\widehat{\beta}_{j\to\mu}\right\}.$ (76)
@@
$\mathbb{E}_{X}[(\mathrm{m}_{i}^{(t)}-\beta_{i}^{(0)})^{2}]\simeq\frac{1}{\alpha^{2}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{j\neq i}^{p}X_{\mu j}\beta_{j}^{(0)}+\xi_{\mu})-\sum_{j\neq i}^{p}X_{\mu i}X_{\mu j}\widehat{\beta}_{j\to\mu}\right\}^{2}$
$\simeq\frac{1}{\alpha}\left\{(\sum_{j\neq i}^{p}X_{\mu j}^{2}(\beta_{j}^{(0)})^{2}+\xi_{\mu}^{2})-2\sum_{k\neq i}X_{\mu k}^{2}\beta_{k}^{(0)}\widehat{\beta}_{k\to\mu}+\sum_{j\neq i}^{p}X_{\mu j}^{2}(\widehat{\beta}_{j\to\mu})^{2}\right\}$
$\simeq\frac{1}{\alpha}\left\{(\frac{1}{p}\sum_{j\neq i}^{p}(\beta_{j}^{(0)})^{2}+\sigma_{\xi}^{2})-\frac{2}{p}\sum_{k\neq i}\beta_{k}^{(0)}\widehat{\beta}_{k\to\mu}+\frac{1}{p}\sum_{j\neq i}^{p}(\widehat{\beta}_{j\to\mu})^{2}\right\}$
$\simeq\frac{1}{\alpha}\left\{\frac{1}{p}\sum_{j\neq i}\left(\beta_{j}^{(0)}-\widehat{\beta}_{j}\right)^{2}+\sigma_{\xi}^{2}\right\}=\frac{E^{(t)}}{\alpha},$ (77)

### 03 Apr 2026 v2

@@ 1 Introduction
A common approach to achieve differential privacy is to randomize the process of parameter estimation via a randomized mechanism. Several such mechanisms have been proposed, including output perturbation, where noise is added directly to the estimated parameters (Dwork et al., 2006), and objective perturbation, where noise is injected into the loss function before optimization (Chaudhuri et al., 2011). A fundamental challenge in all such approaches is balancing the privacy-utility trade-off (Dwork et al., 2014): stronger privacy guarantees typically require more randomness, which can deteriorate statistical accuracy and model interpretability.
In sparse estimation problems such as LASSO, output perturbation can destroy sparsity patterns and deteriorate statistical efficiency (Chaudhuri and Monteleoni, 2008). A natural approach is to first estimate the support of the parameters and then apply perturbation only to the selected components. However, such two-stage procedures require privatizing the support selection step itself, which consumes additional privacy budget by the composition property of differential privacy (Dwork et al., 2014). In contrast, objective perturbation implicitly accounts for the regularization structure of the estimator, and can better preserve sparsity without an explicit support selection step. Despite this appeal, existing theoretical analyses of perturbation-based mechanisms face fundamental limitations in high-dimensional sparse settings; see Section 1.1 for a detailed discussion.
@@
To better understand these interactions, we focus on the high-dimensional regime where both the number of samples and parameters grow to infinity with a fixed ratio $n/p\to\alpha\ll\infty$. In such settings, worst-case DP guarantees can be overly conservative, as they are dominated by rare pathological configurations that do not reflect typical data realizations. We instead adopt a typical-case perspective, analyzing privacy leakage under random design and noise via the on-average KL divergence (OnAveKL) (Wang et al., 2016), which admits an operational interpretation in terms of hypothesis testing: a small OnAveKL implies that even an optimal adversary cannot reliably infer the presence of an individual data point.
@@
Building on this formulation, we analyze the high-dimensional LASSO under both output perturbation and objective perturbation using approximate message passing (AMP). AMP provides an asymptotic characterization of high-dimensional estimators under random design, enabling a typical-case analysis of generalization error and privacy leakage jointly, without relying on restricted eigenvalue assumptions or worst-case sensitivity bounds.
@@ 1.1 Related Works
We review related work from the perspectives of private estimation, empirical risk minimization, and high-dimensional asymptotic analysis, with a focus on the gap between worst-case guarantees and typical-case behavior in high-dimensional regimes.
Perturbation-based mechanisms for private estimation.
Several works have developed perturbation-based mechanisms for differentially private statistical estimation (Chaudhuri et al., 2011; Smith, 2011). However, existing analyses rely on assumptions incompatible with sparse estimation: strong convexity excludes $\ell_{1}$-type penalties (Chaudhuri et al., 2011), and asymptotic normality fails for sparse estimators (Smith, 2011). Kifer et al. (2012) extended objective perturbation to non-differentiable regularizers via a two-stage approach, though their analysis requires restricted strong convexity. Talwar et al. (2015) achieved nearly optimal excess risk of $\tilde{O}(1/n^{2/3})$ for private LASSO without such structural assumptions, but these results are established in a finite-sample regime and do not characterize behavior in the proportional high-dimensional limit $p/n\to\delta>0$, nor the interplay between privacy noise, regularization, and generalization performance.
Private empirical risk minimization.
In the broader context of private empirical risk minimization, Bassily et al. (2014) established minimax optimal rates for convex ERM under only Lipschitz and boundedness assumptions, providing matching upper and lower bounds for both $(\epsilon,0)$- and $(\epsilon,\delta)$-differential privacy. While these guarantees are foundational, they are worst-case in nature and do not characterize behavior in the proportional high-dimensional limit. Related developments include methods that exploit additional structure to mitigate the dimensional dependence inherent in private optimization. For example, Wang et al. (2018) shows that, under smoothness or model assumptions, the sample complexity of ERM under local differential privacy can be significantly improved via polynomial approximation. Similarly, Zhou et al. (2020) proposes a differentially private SGD algorithm that leverages low-dimensional gradient structure to reduce the dependence on the ambient dimension. While these approaches rely on additional structural assumptions to improve upon worst-case bounds, their analyses are still primarily framed in worst-case terms and do not capture the typical-case behavior in high-dimensional regimes.
On-average privacy measures.
On-average privacy measures such as OnAveKL (Wang et al., 2016) have been proposed to capture typical-case privacy leakage, admitting an operational interpretation in terms of hypothesis testing and membership inference. However, existing studies have largely focused on low-dimensional or unregularized settings, leaving open the question of how high dimensionality, sparsity, and regularization jointly affect on-average privacy, particularly in regimes where precise asymptotic characterization becomes possible.
High-dimensional analysis via AMP.
Approximate message passing (AMP) provides a powerful framework for the precise asymptotic characterization of high-dimensional estimators under random design through state evolution (Donoho et al., 2009; Bayati and Montanari, 2011). It has been successfully applied to analyze the typical-case performance of regularized estimators such as the LASSO in the proportional high-dimensional regime. However, its application to privacy-constrained estimation problems remains limited. In this work, we bridge this gap by combining AMP with OnAveKL, enabling a typical-case analysis that jointly captures the effects of privacy mechanisms and statistical estimation in high-dimensional sparse regression, and quantitatively characterizes the trade-off between privacy and generalization.
@@ 1.2 Our contribution
  • Non-monotonic effects of objective noise on privacy and stability: Our analytical characterization provides practical guidance for selecting the noise level to balance prediction accuracy and privacy. In particular, we identify regimes where increasing noise leads to non-monotonic effects on stability and privacy under objective perturbation.
Taken together, our results provide a quantitative characterization of privacy–accuracy trade-offs in high-dimensional sparse estimation, complementing existing worst-case analyses.
@@ 7 Conclusion
$\mathrm{m}_{i}^{(t)}$ $={\Sigma}_{i}^{(t)}\sum_{\mu=1}^{n}\left\{X_{\mu i}\frac{y_{\mu}-\widehat{y}_{\mu}^{\setminus}}{1+{s_{\theta}}}+\frac{X_{\mu i}^{2}\widehat{\beta}_{i}^{(t)}}{1+{s_{\theta}}}\right\}$ (72)
$\frac{1}{{\Sigma}_{i}^{(t)}}$ $=\sum_{\mu=1}^{n}\frac{X_{\mu i}^{2}}{1+{s_{\theta}}}.$ (73)
@@
$=\frac{1}{\sum_{\mu=1}^{n}X^{2}_{\mu i}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{i=1}^{p}X_{\mu i}\beta_{i}^{(0)}+\xi_{\mu}-\sum_{i=1}^{p}X_{\mu i}\widehat{\beta}_{i\to\mu}^{(t)})+X_{\mu i}^{2}\widehat{\beta}_{i\to\mu}^{(t)}\right\}$
$=\frac{1}{\sum_{\mu=1}^{n}X^{2}_{\mu i}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{i=1}^{p}X_{\mu i}\beta_{i}^{(0)}+\xi_{\mu})-\sum_{j\neq i}^{p}X_{\mu i}X_{\mu j}\widehat{\beta}_{j\to\mu}^{(t)}\right\}$ (74)
@@
$\mathrm{m}_{i}^{(t)}-\beta_{i}^{(0)}$ $=\frac{1}{\sum_{\mu=1}^{n}X^{2}_{\mu i}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{j=1}^{p}X_{\mu j}\beta_{j}^{(0)}+\xi_{\mu})-\sum_{j\neq i}^{p}X_{\mu i}X_{\mu j}\widehat{\beta}_{j\to\mu}^{(t)}-X_{\mu i}^{2}\beta_{i}^{(0)}\right\}$
$=\frac{1}{\sum_{\mu=1}^{n}X^{2}_{\mu i}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{j\neq i}^{p}X_{\mu j}\beta_{j}^{(0)}+\xi_{\mu})-\sum_{j\neq i}^{p}X_{\mu i}X_{\mu j}\widehat{\beta}_{j\to\mu}^{(t)}\right\}.$ (76)
@@
$\mathbb{E}_{X}[(\mathrm{m}_{i}^{(t)}-\beta_{i}^{(0)})^{2}]\simeq\frac{1}{\alpha^{2}}\sum_{\mu=1}^{n}\left\{X_{\mu i}(\sum_{j\neq i}^{p}X_{\mu j}\beta_{j}^{(0)}+\xi_{\mu})-\sum_{j\neq i}^{p}X_{\mu i}X_{\mu j}\widehat{\beta}_{j\to\mu}^{(t)}\right\}^{2}$
$\simeq\frac{1}{\alpha}\left\{(\sum_{j\neq i}^{p}X_{\mu j}^{2}(\beta_{j}^{(0)})^{2}+\xi_{\mu}^{2})-2\sum_{k\neq i}X_{\mu k}^{2}\beta_{k}^{(0)}\widehat{\beta}_{k\to\mu}^{(t)}+\sum_{j\neq i}^{p}X_{\mu j}^{2}(\widehat{\beta}_{j\to\mu}^{(t)})^{2}\right\}$
$\simeq\frac{1}{\alpha}\left\{(\frac{1}{p}\sum_{j\neq i}^{p}(\beta_{j}^{(0)})^{2}+\sigma_{\xi}^{2})-\frac{2}{p}\sum_{k\neq i}\beta_{k}^{(0)}\widehat{\beta}_{k\to\mu}^{(t)}+\frac{1}{p}\sum_{j\neq i}^{p}(\widehat{\beta}_{j\to\mu}^{(t)})^{2}\right\}$
$\simeq\frac{1}{\alpha}\left\{\frac{1}{p}\sum_{j\neq i}\left(\beta_{j}^{(0)}-\widehat{\beta}_{j}^{(t)}\right)^{2}+\sigma_{\xi}^{2}\right\}=\frac{E^{(t)}}{\alpha},$ (77)